### v2

# ECCO: Can We Improve Model-Generated Code Efficiency Without Sacrificing Functional Correctness?

Abstract
Although large language models (LLMs) have been largely successful in generating functionally correct programs, conditioning models to produce efficient solutions while ensuring correctness remains a challenge. Further, unreliability in benchmarking code efficiency is a hurdle across varying hardware specifications for popular interpreted languages such as Python. In this paper, we present ECCO, a reproducible benchmark for evaluating program efficiency via two paradigms: natural language (NL) based code generation and history-based code editing. On ECCO, we adapt and thoroughly investigate the three most promising existing LLM-based approaches: in-context learning, iterative refinement with execution or NL feedback, and fine-tuning conditioned on execution and editing history. While most methods degrade functional correctness and moderately increase program efficiency, we find that adding execution information often helps maintain functional correctness, and NL feedback enhances more on efficiency. We release our benchmark to support future work on LLM-based generation of efficient code. [^1] https://github.com/CodeEff/ECCO
[…]
ECCO: Can We Improve Model-Generated Code Efficiency
Without Sacrificing Functional Correctness?
Siddhant Waghjalethanks:  Equal contribution, ordered by a coin toss.     Vishruth Veerendranath11footnotemark: 1     Zora Zhiruo Wang     Daniel Fried Language Technologies Institute, Carnegie Mellon University {swaghjal, vveerend, zhiruow, dfried}@cs.cmu.edu
1 Introduction
The ability to write efficient code is a cornerstone of software development (Li et al., 2022). While large language models (LLMs) have shown remarkable progress in generating functionally correct code (Roziere et al., 2023; Guo et al., 2024), the ability to generate solutions that are both correct and efficient remains elusive (Shypula et al., 2021, 2024).
Current methods for optimizing program efficiency improve performance measured by execution time. However, this apparent success often comes at the cost of severely decreasing the functional correctness (Shypula et al., 2024). An example of this issue is illustrated in Figure 1: When optimizing the program on the left, models sometimes perform spurious optimizations that, although they reduce the runtime, make the program no longer functionally correct so that it fails all test cases. On the other hand, a correct optimization (bottom right) — that improves efficiency while maintaining functional correctness — is often harder to achieve for current LMs. This spurious optimization is certainly undesirable in practice, and can even increase debugging time for software developers (Li et al., 2022; Cummins et al., 2023a). To achieve the goal of real and robust program optimization, we ask: Can LMs improve program efficiency without sacrificing functional correctness?
Figure 1: Correctness-preserving versus spurious optimization when optimizing a linear search algorithm with binary search on a sorted list. Spurious optimization can reduce runtime, but add errors that cause the program to be incorrect. In contrast, a true optimization reduces runtime and remains correct, as we emphasized in ECCO.
[…]
Further, to perform reliable and reproducible executions, we introduce an evaluation setup using a cloud-hosted code execution engine, Judge0 Došilović and Mekterović (2020), which produces stable execution output on correctness, runtime, and memory usage, thanks to its agnostic nature to local hardware specifications. It supports up to 66 programming languages (PLs), allowing future work to extend to other languages.
To explore various correctness-preserving program optimization methods, we evaluate three classes of methods on ECCO— in-context learning, iterative refinement, and fine-tuning, across a suite of open-source language models pre-trained on code — which previous works have used to attempt to improve efficiency, while however overlooking their effects on correctness. We find that execution information and fine-tuning help maintain functional correctness, and NL-involved prompting often yields higher efficiency improvements. However, we broadly reconfirm findings that no existing methods can improve time/space efficiency without sacrificing functional correctness. We hope ECCO can serve as a solid testbed for program optimization, and call for more efforts in advancing correctness-preserving program optimizations.
2 Related Work
Benchmarks for Code Efficiency
Some works have proposed benchmarks for optimizing program assembly code (Bunel et al., 2016; Shypula et al., 2021; Shi and Zhang, 2020; Cummins et al., 2023b). More recently, Shypula et al. (2024) target C++ program speedups, and Huang et al. (2024) evaluate the efficiency of Python solutions for LeetCode coding interview problems (Niu et al., 2024). Although most efforts on LLM-based code generation focus on evaluating functional correctness (e.g., Chen et al. 2021), some works evaluate code efficiency (Moudgalya et al., 2023; Sikka et al., 2020; Jeon et al., 2023; Baik et al., 2024) by classifying the time complexity of programs. However, these works are limited in their single-reference evaluation paradigm, assembly language support, or by the limited problem space on LeetCode (Coignion et al., 2024). Our work supports reliable evaluation across arbitrary coding problems and the widely-used Python language.
Evaluating Program Efficiency
It is challenging to robustly evaluate program efficiency, due to varying hardware platforms and setups. Previous works have evaluated the efficiency of code by executing code in a local software environment Singhal et al. (2024); Huang et al. (2024) or using containerized environments on local hardware (Khan et al., 2023), but this can result in varying runtime and memory usage across hardware, thus causing irreproducible evaluations. An alternative approach is to use an architecture simulator Shypula et al. (2024) which ensures the execution of each program is exactly simulated at the hardware level, but is limited to compiled languages such as C++. A hardware counter though reliable Liu et al. (2024), cannot measure memory usage. For popular interpreted languages such as Python and Java, some use LeetCode’s execution engine (Niu et al., 2024; Coignion et al., 2024), but with a restricted space of testable problems. In our work, we propose an evaluation setup using an accessible cloud computing instance that ensures consistent virtual hardware and reliable benchmarking.
Program Optimization Approaches
To start, some works explore in-context learning to optimize program efficiency (Huang et al., 2024), with retrieval methods to select relevant examples (Gao et al., 2024; Shypula et al., 2024). Beyond vanilla prompting, iterative prompting methods (Madaan et al., 2024; Shinn et al., 2024; Ridnik et al., 2024) have been explored to improve specific aspects of generation, by incorporating feedback from an LM or external modules. Meanwhile, finetuning has been proposed with self-play, synthetic preferences and problem-oriented data Shypula et al. (2024); Gee et al. (2024); Ye et al. (2024). However, none of these methods have been rigorously studied for correctness-preserving optimization. We fill in this gap and provide systematic studies of all methods.
3 The ECCO Benchmark
In this section, we first introduce our evaluation platform (§3.1), then describe the construction process of our ECCO benchmark (§3.2), and lastly, introduce our two task formulations with corresponding evaluation metrics (§3.3).
3.1 Evaluation Platform
To reliably evaluate program efficiency in both runtime and memory usage, we need to first establish a robust and reproducible evaluation platform. However, evaluating program efficiency is challenging, as resource usage statistics vary greatly across hardware and setups (Singhal et al., 2024; Huang et al., 2024).
To ensure stable execution for interpreted languages such as Python, we propose to use a reproducible cloud compute instance that ensures the same virtual hardware, as illustrated in Figure 2. Specifically, we use an EC2 instance (detailed in §B), and execute the code within a code execution engine Judge0 (Došilović and Mekterović, 2020). Note that our recipe can easily extend to over 60 programming languages that the Judge0 engine supports. This is set up as a sandboxed Docker container within the instance, which thus ensures an isolated setup for secure and reproducible code execution. Our platform is similar to evaluating generated code on LeetCode execution console Niu et al. (2024); Coignion et al. (2024), but applies to arbitrary coding problems and is not limited to questions available on LeetCode. [^2] We make the Amazon Machine Image (AMI) for the setup available to enable reproducible benchmarks. We leave evaluations in other languages for future work.
Figure 2: Evaluation platform using Judge0.
3.2 Benchmark Construction
Our goal is to collect programming problems, each with an NL description, and multiple functionally-correct solutions at varied efficiency levels.
Problem Selection
We collect programming problems from the IBM CodeNet dataset (Puri et al., 2021), which contains competitive programming problems with NL descriptions, user program submissions, and other metadata, scraped from the AIZU and AtCoder online judging systems. CodeNet problems mostly require algorithmic techniques such as data structure optimization.
[…]
Next, we split the pairs and group programs by their associated problem ID. We filter out all problems with less than two solutions to ensure that each NL problem description has multiple associated solutions, to enable program optimization based on code editing history. We then remove the programs that are repetitive or cannot successfully execute due to syntax errors or test case failures.
Our curated dataset was partitioned into three subsets: train, validation, and test, with each split consisting of a distinct set of problems. In the end, the process yields 1,380 unique problems in ECCO in total.
Test Case Collection
To evaluate functional correctness, we require test cases. We collect (i) the original test cases for each problem from CodeNet, and (ii) the additional tests from the AlphaCode project Li et al. (2022). Each test case contains the program inputs as well as expected outputs when executing canonical program solutions on these inputs. With these two sets of test cases, we simulate a realistic coding setting where one can refer to (i) as the public test cases for debugging or other accuracy-improving purposes, $T_{public}$, and (ii) as private test cases to conduct final execution-based evaluations on the programs, $T_{private}$.
3.3 Task Formulation and Evaluation
We propose two formulations for the program optimization task, namely NL-instructed generation and history-based program editing. In this section, we introduce the data we use for each formulation, and our evaluations of program correctness, runtime, and memory usage.
3.3.1 History-Based Program Editing
Our first paradigm follows previous work on program optimization Shypula et al. (2024), where we facilitate a history-based editing paradigm. Concretely, we give a previous, presumably slow, version of the solution program, $p_{in}$. We then prompt LMs to edit the code to generate a more efficient version $p_{out}$, denoted as $CodeLM(p_{in})\rightarrow p_{out}$, where $p_{out}$ is expected to run faster than $p_{in}$.
Evaluating Speedup and Memory Reduction
Using the (slow, fast) program pairs remaining after post-processing in §3.2, we evaluate the relative speedup and memory reduction of the model-generated program against the input program on private test cases $T_{private}$.
We adopt the speedup metric introduced by Shypula et al. (2024), which is formulated as:
$Speedup=\frac{\text{Runtime of $p_{in}$}}{\text{Runtime of $p_{out}$}}$ (1)
Similarly, to evaluate improvement in memory usage, we introduce a memory reduction metric as:
$Memory\ Reduction=\frac{\text{Memory of $p_{in}$}}{\text{Memory of $p_{out}$}}$ (2)
Figure 3: Illustration of history-based editing.
3.3.2 NL-Instructed Generation
In addition, we support the most common NL-to-code generation setup: given the NL description $d$ of a problem, we ask the LM to generate the program solution $p$, as $CodeLM(d)\rightarrow p$. Our goal is for the code LM to generate an efficient and correct solution $p$. We execute $p$ on the private test cases $T_{private}$ to evaluate its performance.
Figure 4: Illustration of NL-instructed generation.
Solution Program Spectrum
To evaluate relative runtime and memory efficiency, we measure where a model-generated program lies on the spectrum of all user-submitted programs to that problem. We use the Judge0 evaluation platform (§3.1) to measure the runtime and memory usage.
Evaluating Percentile over the Spectrum
We introduce runtime and memory percentile to measure the efficiency of the model-generated program over the solution spectrum for a given problem as:
$Runtime\ \%=\frac{\text{\# Slower user programs}}{\text{Total \# of user programs}}$ (3)
$Memory\ \%=\frac{\text{\# Programs w/ more memory}}{\text{Total \# of user programs}}$ (4)
3.3.3 Evaluating Functional Correctness
To measure if program correctness is preserved, a key metric is the functional correctness of model-generated programs. We adopt the pass@1 metric introduced by Chen et al. (2021), which samples one program from the model and measures whether the generated program passes all test cases.
3.4 ECCO Feature Analysis
[…]
Split # Problems # Pairs # Avg. Test Cases
Public Private
Train 1262 48386 3.14 17.21
Val     69 2359 3.17 17.25
Test     48 794 3.29 20.00
Table 1: ECCO dataset statistics.
Figure 5: Iterative refinement methods utilizing different forms of feedback. Self-Refine uses Natural Language feedback, Exec-Refine uses raw execution results on $T_{public}$ and NL+Exec-Refine uses NL reflection of execution.
4 Efficiency-Improving Approaches
We explore various top-performing code generation approaches to improve program efficiency, while maintaining functional correctness, including in-context learning (§4.1), iterative refinement (§4.2), and fine-tuning (§4.3).
4.1 In-Context Learning
We explore two mainstream prompting strategies: instruction prompting and few-shot learning.
Instruction prompting
Many LMs perform better when incorporating instructions Ouyang et al. (2022); Wei et al. (2022). We use two prompts: $I_{gen}$ for NL-based generation which instructs models to generate correct and efficient programs; and $I_{eff}$ for history-based editing which instructs models to optimize the input program. $I_{eff}$ is adapted from PIE Shypula et al. (2024) and Self-Refine Madaan et al. (2024). See §D for details.
Few-Shot Learning
We add few-shot example demonstrations (Brown et al., 2020): for the NL-based setting, using (NL, fastest program) pairs; for history-based editing, using (slow program, fast program) pairs. We randomly sample examples from the train set as the few-shot examples.
4.2 Iterative Refinement
We explore three methods (illustrated in Figure 5 ) to iteratively refine the generated code to be more efficient, which intuitively aligns with the way that humans improve code (Madaan et al., 2024).
Self-Refine with NL Feedback
We adopt self-refine (Madaan et al., 2024) that prompts LMs to iteratively examine the output and refine it. More concretely, (1) we first prompt the LM to generate a candidate solution; (2) we ask the same model to produce NL reasoning about why the code is incorrect and/or inefficient; and (3) we input the original input and the feedback from (2) to the model and ask it to generate an updated solution.
Exec-Refine with Interpreter Feedback
We propose an alternative refinement strategy that obtains deterministic execution feedback from the interpreter, by running the program over $T_{public}$. [^5] Models do not have access to the private test cases we finally evaluate on. If test cases are passed, the execution result provides the runtime and memory information; otherwise, this feedback provides interpreter error logs. Both correctness and efficiency can be informed via this feedbackare.
NL+Exec Refine: NL Feedback on Interpreter Results
To allow feedback both in the forms of NL and execution outputs, we ground the LM feedback on execution results, inspired by the Reflexion feedback paradigm Shinn et al. (2024). Specifically, we first obtain the execution results as in exec-refine, then ask the LM to write NL feedback on the incorrect/inefficient parts in the code, and use this as additional input in the refinement turn.
Model Setting History-based Editing NL-instructed Generation
pass@1 speedup memory reduction pass@1 runtimememory%
StarCoder2 instruct 49.4 1.49 1.24   4.2 50.64 55.72
few-shot 49.8 1.70 1.07   2.1 11.40 50.17
CodeGemma instruct 42.5 1.43 1.10 18.8 41.70 51.83
few-shot 43.9 1.07 1.06 22.9 62.80 67.33
WizardCoder instruct 34.2 1.58 1.18 14.6 54.29 84.53
few-shot 27.4 1.38 1.12 14.6 58.69 71.00
CodeLLaMa instruct 57.5 1.44 1.11   8.3 45.30 74.18
few-shot 22.5 1.63 1.26   8.3 42.66 67.21
DeepseekCoder instruct 29.8 2.11 1.28 18.8 59.01 75.86
few-shot 35.2 2.26 1.20 22.9 55.52 66.09
GPT-4o instruct 66.6 1.64 1.10 52.1 46.01 59.21
few-shot 65.8 1.62 1.12 41.7 49.87 64.44
Table 2: Results using In-Context Learning approaches (instruction-prompting and few-shot learning)
4.3 Fine-tuning
We also explore three fine-tuning methods beyond prompting-alone approaches.
Vanilla Fine-tuning
In this vanilla training setting, we leverage (NL, program) pairs and (slow program, fast program) pairs to train models independently for each paradigm. We format the data for both similarly to the in-context learning prompts (§4.1), and finetune on a causal language modelling task on the formatted data for each of the two paradigms independently.
Execution Conditioned Fine-tuning
Beyond fine-tuning with basic contexts, we posit that further conditioning on execution results could help. Therefore, we include execution results of PASS/FAIL status, runtime, and memory usage for each public test case for the input program.
Trajectory Conditioned Fine-tuning
For history-based editing, we further propose trajectory-conditioned fine-tuning, by adding a trajectory history of programs written by the same user for the given problem in context. We first collect all problems with at least three programs submitted by the same user, and treat the series of programs as a trajectory. From each qualified trajectory, we designate the fastest code as the target output, and sample three other intermediate programs at the 0th, 33rd, and 66th percentile steps to use as inputs. We aim to allow the model to learn from the step-by-step improvements that led to the optimal solution, capturing the problem-solving process in addition to just the inputs and targets.
5 Experiments
5.1 Experimental Setup
Models
We experiment with several best-performing LMs pre-trained on code. Specifically, we evaluate CodeLlama-13B Roziere et al. (2023), DeepSeekCoder-7B-v1.5 Guo et al. (2024), CodeGemma-7B Team et al. (2024), WizardCoder-13B-Python Luo et al. (2023), StarCoder2-15B Lozhkov et al. (2024). We use the instruction-tuned versions of all of these open-checkpoint models unless indicated otherwise. We also use the proprietary GPT-4o model for no-training methods.
5.2 Results and Analysis
5.2.1 In-Context Learning
As shown in Table 2, all methods either reduce pass@1 of the program by a large margin (in editing mode) or obtain low pass@1 (generation mode). Comparing the two paradigms, history-based editing results in a substantially higher pass@1 by referring to a base correct program, compared to NL-instructed generation which lacks a base program to start from. GPT-4o obtains a much higher pass@1 than all models in both paradigms.
History-based editing
While in-context learning can effectively speed up the input program by 7777126126126126%, but compromises correctness, dropping it to 22.522.522.522.566.666.666.666.6, and uses more memory. Few-shot shows this trend more explicitly than instruct. Besides the limitations of LMs, this may be caused by the sampled few-shot demonstrations being algorithmically less relevant to the problem at hand.
NL-instructed generation
While DeepseekCoder and CodeGemma’s pass@1 improves by $4.1\%$ with few-shot, GPT-4o and StarCoder2’s pass@1 drops by $2.1-10.4\%$. Similarly for the efficiency metrics, CodeGemma and GPT-4o see an improvement whereas other models do not. GPT-4o significantly outperforms other models at pass@1 by 2.2$\times$, however there is no clear winner for efficiency. This highlights the complex trade-offs between correctness and efficiency specifically in the NL-instructed generation task.
Model Setting History-based Editing NL-instructed Generation
pass@1 speedup memory reduction pass@1 runtimememory%
StarCoder2 pre-refine 49.4 1.49 1.24   4.2 50.64 55.72
self-refine 26.7 1.55 1.30   2.1   5.79 71.50
exec-refine 39.5 1.49 1.23   2.1 29.27 55.79
nl+exec refine 26.1 2.13 1.26   2.1 5.79 81.69
CodeGemma pre-refine 42.5 1.43 1.10 18.8 41.70 51.83
self-refine 15.1 2.08 1.15   6.3 41.23 59.18
exec-refine 33.2 1.59 1.12 18.8 39.26 54.81
nl+exec refine 29.8 1.54 1.14 14.6 33.24 38.70
WizardCoder pre-refine 34.2 1.58 1.18 14.6 54.29 84.53
self-refine   8.5 2.16 1.23   8.3 44.86 88.77
exec-refine 20.9 1.60 1.13 12.5 44.12 76.86
nl+exec refine 18.3 2.90 1.30 12.5 31.92 79.19
CodeLLaMa pre-refine 57.5 1.44 1.11   8.3 45.30 74.18
self-refine 15.8 2.02 1.22   2.1 32.16 99.42
exec-refine 54.6 1.51 1.12   4.2 44.09 85.28
nl+exec refine 16.2 1.37 1.02   4.2 66.00 70.79
DeepseekCoder pre-refine 29.8 2.11 1.28 18.8 59.01 75.86
self-refine 13.6 2.73 1.35   8.3 29.65 65.26
exec-refine 27.4 2.34 1.24 20.8 55.08 73.78
nl+exec refine 19.6 3.54 1.37 14.6 49.08 85.15
GPT-4o pre-refine 66.6 1.64 1.10 52.1 46.01 59.21
self-refine 47.8 2.72 1.25 37.5 51.12 44.03
exec-refine 60.8 2.19 1.22 52.1 52.55 59.47
nl+exec refine 58.8 2.39 1.22 47.9 49.79 47.46
Table 3: Results with iterative refinement approaches. Feedback in NL (self & nl+exec) improves efficiency better, whereas raw execution feedback (exec) maintains correctness more effectively.
5.2.2 Iterative Refinement
Table 3 shows all results with iterative refinement methods. As a reference for the refinement approaches, we measure the LM-generated code in the first attempt at optimization without any refinement, and denote this method as pre-refine.
History-based editing paradigm
While all methods can effectively speed up the program, methods that involve NL feedback (self-refine and nl+exec refine) achieve the highest speedup across models. exec-refine consistently yields the highest pass@1 for all models, by 3.43.43.43.438.838.838.838.8 points more than the other two methods. We conjecture that execution outputs are better representations to inform functional correctness than NL descriptions. Although it is easier to convey high-level optimization strategies in NL, conveying the functional correctness is harder. Overall, although the models are instructed to emphasize both correctness and efficiency, there seems to be an implicit trade-off between them. Additional analysis is in §A.
NL-instructed generations
We observe similar patterns as the editing mode, that exec-refine best maintains functional correctness, and two other NL-involved approaches improve runtime/memory efficiency. Compared to the in-context learning results in Table 2, iterative refinement significantly improved memoryfor all the models, with the best method showing an average improvement of $12.06\%$ over the instruct method. However, the impact on runtimeshows varying results among the different models.
(a) Pass@1
(b) Speedup
(c) Memory Reduction
Figure 6: Performance of DeepseekCoder over multiple iterations of refinement. The improvement in efficiency is outweighed by the consistent drop in pass@1.
5.2.3 Fine-tuning
We perform parameter-efficient fine-tuning on CodeLLaMa-7B and DeepseekCoder-7B, the best-performing classes of models on the correctness and efficiency metrics in our prompting experiments respectively.
History-based editing
As shown in Table 4, fine-tuning is the most effective method in maintaining correctness in the editing paradigm. Especially for DeepseekCoder, compared to the highest prompting results $35.2$ using few-shot examples, vanilla and execution-conditioned tuning improves by $6.9$ and $7.8$ points, and trajectory-conditioned tuning further gains a $34.6$ point increase overall. This suggests that adding user-specific coding trajectories can help ground models into the optimization mode and substantially improve output correctness.
Model Method pass@1 speedup mem.red.
CodeLLaMa-7B Vanilla 43.0 1.11 1.01
Execution 45.0 1.41 1.04
Trajectory 70.2 1.01 1.00
DeepseekCoder Vanilla 42.1 1.11 1.01
Execution 43.0 1.16 1.02
Trajectory 69.8 1.01 1.00
Table 4: Fine-tuning results for history-based editing.
NL-instructed generation
In the more complex NL-instructed generation task shown in Table 5, fine-tuning is effective in improving the efficiency for CodeLLaMa but not for DeepseekCoder, highlighting the need for more robust fine-tuning methods that can handle trade-offs and effectively maintain correctness in this setting.
However, fine-tuning results in a much less efficiency improvement than prompting-based methods for both paradigms, possibly due to the limited power of parameter efficient fine-tuning.
Model Method pass@1 runtimemem%
CodeLLaMa-7B Pre-trained 8.3 36.74 44.15
Finetuned 8.3 46.81 71.32
DeepseekCoder Pre-trained 18.8 59.01 75.86
Finetuned 16.7 43.61 67.31
Table 5: Fine-tuning results for NL-based generation.
6 Additional Analysis
Does instruction tuning help in-context learning?
As instruct model versions are expected to be better at in-context learning than their base counterparts, we compare base and instruct model versions in the editing paradigm. In Table 6, the base versions get much higher correctness albeit at lower efficiency, showing that base and instruct versions lie at different points on the correctness-efficiency trade-off. As we emphasize both correctness and efficiency aspects in the NL instruction, we conjecture the instruct models take more hints from the input format and emphasize efficiency, while base models primarily emphasize correctness.
Model Version pass@1 speedup mem.red.
CodeLlama Instruct 22.5 1.63 1.26
Base 46.4 1.02 1.00
DeepseekCoder Instruct 35.2 2.26 1.20
Base 45.4 1.04 1.00
CodeGemma Instruct 43.9 1.07 1.06
Base 48.2 1.01 1.00
StarCoder2 Instruct 49.8 1.70 1.07
Base 41.5 1.03 1.00
Table 6: Comparing base and instruct model versions with in-context learning methods.
Does multi-iteration refinement help?
Multiple refinement iterations may improve results by allowing more turns for models to refine. To verify this, we evaluate the iterative refinement method using 1–4 iterations. While self-refine and exec-refine improve program speedup over iterations (6(b)), all methods continuously degrade the pass@1 to various extents. Exec-refine preserves correctness more effectively in further iterations as well. For memory usage (6(c)), exec-refine consistently reduces memory usage, yet other methods exhibit big fluctuations. In general, more iterations can speed up the program, yet further sacrifice correctness, thus has limited gains in terms of correctness-preserving optimization.
Figure 7: Correctness and efficiency of generated code across model sizes for CodeLLaMa and DeepseekCoder. The top row corresponds to History-Based Editing and the bottom row includes NL-Instructed Generation.
Can iterative prompting fix incorrect solutions?
To study whether models can recover from incorrect starting solutions in the history-based editing paradigm, we evaluate on a collection of 157 pairs of programs, ECCO-Fix, where the input code is almost correct: one that passed all public test cases but fails a few private test cases. As shown by Table 7, we show that exec-refine can fix incorrect programs to pass all public and private test cases, with access to only the PASS/FAIL status of public test cases. In comparison, self-refine breaks the correctness of more programs.
Aligning with our findings in earlier sections and §A, exec-refine, with execution information in contexts, can encourage models to generate functionally correct programs.
Model Instruct Self-Refine Exec-Refine
StarCoder2 12.1 9.5 8.9
CodeGemma 10.8 3.1 10.8
WizardCoder 20.3 1.9 14.6
CodeLlama 17.1 7.6 36.9
DeepseekCoder 12.7 5.7 15.2
Table 7: Pass@1 of iterative refinement strategies for models on ECCO-Fix, under the editing paradigm.
How does scale of model impact efficiency and correctness?
We perform experiments across model scales for the CodeLLaMa (7B, 13B, 34B, 70B) and DeepseekCoder-v1 (1.3B, 6.7B, 33B) instruction-tuned model families. As seen in Figure 7 and Table 10, for both History-based Editing and NL-Instructed Generation, scaling model size presents a mix of benefits and trade-offs in terms of efficiency and correctness. For History-based Editing, larger CodeLLaMa models (such as 70B) show better speedup but experience diminishing gains in correctness (Pass@1), with a slight decrease in memory reduction. DeepseekCoder follows a similar pattern but consistently underperforms CodeLLaMa in correctness. In NL-Instructed Generation, scaling models significantly improves correctness, as seen with CodeLLaMa 70B and DeepseekCoder 33B, although giving varying results for runtime and memory percentiles.
7 Conclusion
In this paper, we introduce the ECCO benchmark that enables two paradigms for Python program optimization, using Judge0, a language and platform-agnostic execution framework. We find that execution information and fine-tuning help LLMs maintain code correctness, and prompting with natural language often yields higher efficiency gains. However, we broadly reconfirm findings that no existing method can improve efficiency without sacrificing functional correctness. We hope ECCO can serve as a testbed for program optimization, and we call for more efforts in advancing correctness-preserving program optimizations.
Acknowledgements
We thank Aman Madaan for the helpful discussions in the early stage of the project.
Limitations
Our benchmark establishes a solid foundation for rigorously evaluating the ability to generate efficient solutions, however, we also note that there are limitations to our work.
First, ECCO has only included Python problems so far, but our Judge0 evaluation platform and our benchmark curation recipe are fully reproducible and could be extended to other programming languages of interest. Second, our benchmark currently focuses on competitive programming problems. It is also possible to extend our benchmark to other types of programming problems from more real-world software engineering scenarios.
Due to both limitations, our results may not be comprehensive enough to reflect the quality of model-generated programs on the full spectrum of all programming languages and problems. When using ECCO in practice, we recommend the readers examine the model outputs, in addition to the quantitative results produced by our framework.
Appendix A Iterative Refinement Analysis
Following Shypula et al. (2024), we also evaluate runtime and memory efficiency as a percentage of pairs where the generated code $p_{out}$ is faster/uses lesser memory than $p_{in}$, in addition to the speedup and memory reduction metrics described in §3.3.1.
In Table 8, these percentage optimized metrics clearly indicate that exec-refine is the most consistent iterative refinement method as it achieves the best pass@1, % runtime optimized and % memory optimized across all models. Contrasting these results to Speedup and Memory Reduction in Table 3, we note that while natural language feedback (in self-refine and NL+Exec refine) aids in significantly improving the runtime and memory usage for a few cases (while breaking test cases for others), exec-refine improves runtime and memory in more cases albeit to a smaller degree.
Model Setting History-based Editing
pass@1 % Runtime Opt % Mem. Opt.
StarCoder2 self-refine 26.7 22.5 27.0
exec-refine 39.5 25.5 30.3
nl+exec refine 26.1 22.6 27.0
CodeGemma self-refine 15.1 18.2 29.3
exec-refine 33.2 25.5 34.5
nl+exec refine 29.8 21.0 31.3
WizardCoder self-refine 8.5 16.1 24.4
exec-refine 20.9 20.0 27.2
nl+exec refine 18.3 17.1 26.9
CodeLLaMa self-refine 15.8 15.6 24.5
exec-refine 54.6 27.4 39.6
nl+exec refine 16.2 16.8 23.2
DeepseekCoder self-refine 13.6 26.8 35.8
exec-refine 27.4 31.4 40.0
nl+exec refine 19.6 29.7 36.5
GPT-4o self-refine 47.8 42.5 42.3
exec-refine 60.8 48.8 51.1
nl+exec refine 58.8 48.6 48.1
Table 8: Comparing iterative refinement approaches on % Optimization metrics. Exec-Refine is the best performing approach across all models and metrics.
Appendix B Implementation Details
Generation
We use a maximum length of 1024 tokens, and a temperature of $t=0.4$. We provide two examples for all in-context few-shot experiments
Hardware
We run all our experiments on a mix of A6000 and L40 GPUs. Specifically for the prompting iterative prompting and in-context learning approaches, we use 1-2 GPUs of the type and utilize the vLLM library Kwon et al. (2023) primarily for generating programs efficiently. We perform finetuning on 4 A6000 GPUs.
For Judge0 evaluation virtual hardware setup, we use an m7i.large EC2 instance which has 2 vCPU cores of the 4th Generation Intel Xeon Sapphire Rapids, with 8GB of RAM.
Fine-tuning
We adopt parameter-efficient fine-tuning with LoRA (Hu et al., 2022) due to resource limitations and implement this using the HuggingFace Transformers library  [^6] https://github.com/huggingface/transformers. We optimize the finetuning using Deepspeed  [^7] https://github.com/microsoft/DeepSpeed ZeRo stage 2 with a LoRA rank of 8, alpha of 16 and a per-device batch size of 2. We use the AdamW optimizer, with a learning rate of 1e-3 and a warmup of 100 steps. We train the models on the history-based editing task for a single epoch and fine-tune models for 10 epochs on the NL-instructed generation task.
Appendix C Variance Analysis
We measured the variance for two models (DeepseekCoder and CodeLLaMa) by sampling generations three times and evaluating each generation independently on our Judge setup. The results are included below with 95% Confidence Intervals are shown in Table 9.
Pass@1 has no variance across runs in both paradigms due to the low temperature; runtime efficiency metrics (speedup and runtime %) both only have about  1% Confidence Interval (CI). Memory has close to no variance in the history-based editing setting but shows  5% CI in NL-instructed generation.
Model History-based Editing NL-Instructed Generation
Pass@1 Speedup Memory Reduction Pass@1 Runtime (%) Memory (%)
DeepseekCoder 32.87 $\pm$ 0.00 2.112 $\pm$ 0.022 1.097 $\pm$ 0.000166 18.75 $\pm$ 0.00 57.589 $\pm$ 1.012 67.796 $\pm$ 3.106
CodeLLaMa 57.50 $\pm$ 0.00 1.456 $\pm$ 0.0028 1.114 $\pm$ 0.00173 8.30 $\pm$ 0.00 55.663 $\pm$ 0.720 61.029 $\pm$ 3.818
Table 9: Variance of results for History-based Editing and NL-Instructed Generation
Model Size History-based Editing NL-instructed Generation
Pass@1 Speedup Memory Reduction Pass@1 Runtime (%) Memory (%)
CodeLLaMa 7B 58.4 1.04 1.00   2.1 39.73 26.93
13B 57.5 1.44 1.11   8.3 45.30 74.18
34B 51.9 1.31 1.13   8.3 43.23 70.16
70B 55.3 1.73 1.09 12.5 39.15 64.89
DeepseekCoder 1.3B 44.2 1.36 1.13   8.3 46.71 65.62
6.7B 34.8 1.47 1.16 29.2 33.48 76.19
33B 49.1 1.51 1.09 37.5 43.91 75.61
Table 10: Results for History-based Editing and NL-Instructed Generation across different model sizes.
Appendix D Prompt Details
We illustrate and detail all of the prompts used for the experiments in Figures 8-14.
{mdframed}
[ backgroundcolor=black!10, outerlinewidth=2pt, outerlinecolor=blue, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=15pt, innerleftmargin=15pt ]
Optimize the python program below to be functionally equivalent
but run faster and use less memory.
Here are a few examples:
### Program:
{slow_code_example}
### Optimized (Runtime and Space) version of Program above:
{fast_code_example}
### Program:
[src_code]
### Optimized (Runtime and Space) version of Program above:
Figure 8: Prompt for Instruction prompting $I_{eff}$ along with in-context examples
{mdframed}
[ backgroundcolor=black!10, outerlinewidth=2pt, outerlinecolor=blue, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=15pt, innerleftmargin=15pt ]
Your solution was functionally {CORRECT/INCORRECT}
Here are the run time and memory stats of your code for each test case
-- Stats for test case 0 --
Correct: {PASSED/FAILED}
Run time: 0.009 s
Memory: 3352.0 KB
-- Stats for test case 1 --
Correct: {PASSED/FAILED}
Run time: 0.009 s
Memory: 3316.0 KB
Figure 9: Example of Execution Feedback on public test cases used for Exec-Refine and Execution Conditioned Fine-tuning
{mdframed}
[ backgroundcolor=black!10, outerlinewidth=2pt, outerlinecolor=blue, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=15pt, innerleftmargin=15pt ]
Write a python code which is correct and efficient in terms of runtime and memory
usage for the following problem description.
##Problem Name:
{problem_name}
##Problem Description:
{In detail description of the task}
## Sample Inputs:
{input_test_cases}
##Sample Outputs:
{Expected Output}
Figure 10: Prompt for NL-instructed generation $I_{gen}$
{mdframed}
[ backgroundcolor=black!10, outerlinewidth=2pt, outerlinecolor=blue, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=15pt, innerleftmargin=15pt ]
Give feedback in english for why the code solution below is incorrect
or inefficient and how the program can be fixed.
## Candidate solution:
{most recent code attempt}
## Feedback for incorrectness/inefficiency and how it can be improved:
Figure 11: Prompt used for NL-reasoning to get feedback
{mdframed}
[ backgroundcolor=black!10, outerlinewidth=2pt, outerlinecolor=blue, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=15pt, innerleftmargin=15pt ]
Refine the given incorrect or sub-optimal code solution based on the feedback
specified below.
### Candidate solution:
{previous_code_attempt}
### Feedback for incorrectnes/inefficiency and how it can be improved:
{self-feedback / execution-feedback / NL+Exec-Refine}
### Optimized/Corrected solution based on feedback:
Figure 12: Prompt used for refining code in Self-Refine, Exec-Refine and NL+Exec-Refine
{mdframed}
[ backgroundcolor=black!10, outerlinewidth=2pt, outerlinecolor=blue, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=15pt, innerleftmargin=15pt ]
Based on the execution results, reflect on why the code solution
below was incorrect or inefficient and how the program can be fixed.
{generated code solution}
## Execution Results:
{Execution Feedback for the previous attempt}
### Reflection on incorrectnes/inefficiency and how it can be improved:
Figure 13: Prompt used for NL+Exec-Refine to reflect on Execution results
{mdframed}
[ backgroundcolor=black!10, outerlinewidth=2pt, outerlinecolor=blue, roundcorner=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerrightmargin=15pt, innerleftmargin=15pt ]
##1 iteration program:
{slowest program in trajectory}
##2 iteration program:
{33 percentile fastest program in trajectory}
##3 iteration program:
{66 percentile fastest program in trajectory}
### Final iteration program:
Figure 14: Format of Trajectory-Conditioned Fine-tuning data